• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 09:26:01 on November 10, 2009

    Entry number 298748


    keyword=Current status

    Current major issues we have at this moments are:
    1. High Compton background,
    2. HRS magnet cryogen
    3. Left HRS dipole

    For the Compton issue, experts in MCC are gathered today to trouble-shoot
    the problem and try to do the tuning. This probably will take the whole
    day shift and they don't want to be interrupted by any access.

    For the magnet cryogen, the right HRS magnets tripped due to low
    cryogenic level. The techs are looking into, they seriously need access
    to figure out the problem.

    For the left HRS dipole, Jack also needs access to fix it.

    So current plan is that, since the Compton/beam line experts are
    available and have prepared for quite some time, we let them do the
    tuning today, which means the techs can only start to work on the cryo
    and dipole tomorrow. Currently, the right HRS cryo level seems recovered,
    hopefully, we can still take production data on it tonight.

    A copy of this log entry has been emailed to: xiaochao, rom, yqiang